It's nighttime, and you've dropped your goggles into a 3.2-m-deep swimming pool. If you hold a laser pointer 0.90m above the edge of the pool, you can illuminate the goggles if the laser beam enters the water 2.2m from the edge.

How far are the goggles from the edge of the pool?
Answer #1

here ,

depth of pool , d = 3.2 m

Now, let i be the angle of incidence

i = arctan(2.2/.90)

i = 67.8 degree

Using snell's law ,

n1 * sin(i) = n2 * sin(r)

1 * sin(67.8) = 1.33 * sin(r)

r = 44.1 degree

Now,

distance of googles = 2.2 + d*tan(r)

distance of googles = 2.2 + 3.2 * tan(44.1)

distance of googles = 5.3 m

the googles are 5.3 m from the edge